# Neural Combinatorial Optimization for Real-World Routing

### v1

@@ 5.1.1. Initial Embedding Module
The initial embedding module synthesizes distance-related features with node characteristics to create comprehensive node representations. A key aspect of our approach is effectively integrating two complementary spatial features: distance matrix information and coordinate-based relationships. For distance matrix information, we employ a selective sampling strategy that captures the most relevant node relationships while maintaining computational efficiency. Given a distance matrix $\mathbf{D}\in\mathbb{R}^{N\times N}$, we sample $k$ nodes for each node $i$ according to probabilities inversely proportional to their distances:
(4) $p_{ij}=\frac{1/d_{ij}}{\sum_{j=1}^{N}1/d_{ij}}$
@@
(5) $\mathbf{f}_{\text{dist}}=\text{Linear}(\mathbf{d}_{\text{sampled}})$
@@
(6) $\mathbf{f}_{\text{coord}}=\text{Linear}(\mathbf{x}_{\text{coord}})$
@@
(7) $\mathbf{h}=\mathbf{g}\odot\mathbf{f}_{\text{coord}}+(1-\mathbf{g})\odot\mathbf{f}_{\text{dist}}$
@@
(8) $\mathbf{g}=\sigma(\text{MLP}([\mathbf{f}_{\text{coord}};\mathbf{f}_{\text{dist}}]))$
@@
(9) $\mathbf{h}^{r}_{\text{comb}}=\text{MLP}([\mathbf{h}^{r};\mathbf{f}_{\text{node}}])$
(10) $\mathbf{h}^{c}_{\text{comb}}=\text{MLP}([\mathbf{h}^{c};\mathbf{f}_{\text{node}}])$
@@
where $\mathbf{f}_{\text{node}}$ represents additional node features such as demand or time windows, which are transformed by an additional linear layer. This dual embedding approach allows the RRNCO model to better capture and process asymmetric relationships (that can include both distances and durations) in real-world routing scenarios.
@@
Table 1. Performance comparison across real-world routing tasks and distributions. We report costs, time, and gaps calculated with respect to best-known solutions (*) from traditional solvers. Horizontal lines separate traditional solvers, node-only encoding learning methods, node and edge encoding, and our RRNCO, respectively, for each task. The lower is better ()($\downarrow$)( ↓ ).
@@
RRNCO 122.693 3.928 35s 123.249 3.996 35s 41.077 4.647 35s
5.1.2. Neural Adaptive Bias for AAFM
Building upon the comprehensive initial embedding module, RRNCO leverages the Adaption Attention-Free Module (AAFM)(Zhou et al., 2024b), which incorporates Attention Free Transformer (Zhai et al., 2021) operations to enhance the model’s ability to capture intricate node relationships through a Neural Adaptive Bias (NAB) mechanism.
The Adaption Attention Free Module (AAFM) proposed in the (Zhou et al., 2024b) uses the following operation:
(11) $\mathbf{\text{AAFM}}(Q,K,V,A)=\sigma(Q)\odot\frac{\exp(A)\cdot(\exp(K)\odot V)}{\exp(A)\cdot\exp(K)}$
where $Q=\mathbf{W}^{Q}\mathbf{h}^{r}_{\text{comb}}$, $K=\mathbf{W}^{K}\mathbf{h}^{c}_{\text{comb}}$, and $V=\mathbf{W}^{V}\mathbf{h}^{c}_{\text{comb}}$. respectively, and $\mathbf{W}^{Q}$, $\mathbf{W}^{K}$, and $\mathbf{W}^{V}$ are learnable weight matrices. $A$ represents an adaptation bias that encodes relationships between nodes. The original paper defines this adaptation bias heuristically as $-\alpha\cdot\log(N)\cdot d_{ij}$, where $\alpha$ is a learnable parameter, $N$ is the number of nodes, and $d_{ij}$ is the distance between nodes $i$ and $j$. Our approach differs in that we do not rely on this heuristic formulation of the adaptation bias. Instead, we propose using a Neural Adaptive Bias (NAB) that learns to capture rich relationships between nodes in a data-driven manner. The NAB takes as input a distance matrix encoding pairwise distances between nodes as well as an angle matrix representing the angular relationships between nodes.
Let $\mathbf{W}_{D},\mathbf{W}_{\Phi}\in\mathbb{R}^{E}$, where $E$ is the embedding dimension:
$\mathbf{D}_{emb}=\text{ReLU}(\mathbf{D}\mathbf{W}_{D})\mathbf{W}^{\prime}_{D}\in\mathbb{R}^{B\times N\times N\times E}$
$\mathbf{\Phi}{emb}=\text{ReLU}(\mathbf{\Phi}\mathbf{W}_{\Phi})\mathbf{W}^{\prime}_{\Phi}\in\mathbb{R}^{B\times N\times N\times E}$
We then apply contextual gating to fuse the distance and angle information. The gate $\mathbf{G}$ is computed as:
$\mathbf{G}=\sigma\left(\left[\mathbf{D}_{emb};\mathbf{\Phi}_{emb}\right]\mathbf{W}_{G}\right)\in\mathbb{R}^{B\times N\times N\times 1}$
where $\left[\mathbf{D}_{emb};\mathbf{\Phi}_{emb}\right]\in\mathbb{R}^{B\times N\times N\times 2E}$ denotes the concatenation of the distance and angle embeddings, $\mathbf{W}_{G}\in\mathbb{R}^{2E}$ is a learnable weight matrix, and $\sigma$ is the sigmoid activation function.
$\mathbf{H}=\mathbf{G}\odot\mathbf{D}_{emb}+(\mathbf{1}-\mathbf{G})\odot\mathbf{\Phi}_{emb}\in\mathbb{R}^{B\times N\times N\times E}$
where $\odot$ denotes element-wise multiplication. Finally, the adaptive bias matrix $\mathbf{A}$ is obtained by projecting the fused embedding $\mathbf{H}$ to a scalar value:
$\mathbf{A}=\mathbf{H}\mathbf{w}_{out}\in\mathbb{R}^{B\times N\times N}$
where $\mathbf{w}_{out}\in\mathbb{R}^{E}$ is a learnable weight vector. The resulting $\mathbf{A}$ matrix serves as an inductive bias that reflects rich node relationships based on both distances and angles, with the contextual gate $\mathbf{G}$ controlling the balance between the two information sources. This Neural Adaptive Bias can then be incorporated into the Attention Attention Free Module (AAFM) operation as follows:
$\text{AAFM}(Q,K,V,\mathbf{A})=\sigma(Q)\odot\frac{\exp(\mathbf{A})\cdot(\exp(K)\odot V)}{\exp(\mathbf{A})\cdot\exp(K)}$
The Neural Adaptive Bias (NAB), applied through the Attention Attention Free Module (AAFM), yields the final node representations $h^{r}_{\textit{F}}$ and $h^{c}_{\textit{F}}$ after passing through the AAFM $l$ times. These representations are the product of RRNCO’s encoding process, leveraging distance and angle embeddings to capture and encode rich, contextually relevant information about the underlying graph structure.
5.2. Decoder
5.2.1. Decoder Architecture
The decoder architecture combines key elements from the ReLD (Huang et al., 2025) and MatNet (Kwon et al., 2021) to effectively process the dense node embeddings generated by the encoder and construct solutions for vehicle routing problems. At each decoding step $t$, the decoder takes as input the row and column node embeddings ($h^{r}_{\textit{F}}$, $h^{c}_{\textit{F}}$) produced by the encoder and a context vector $h_{c}=[h^{r}_{a_{t-1}},D^{t}]\in\mathbb{R}^{d_{h}+d_{attr}}$, where $D^{t}\in\mathbb{R}^{d_{attr}}$ represents dynamic features that capture the state variable $s^{t}$. To aggregate information from the node embeddings, the decoder applies a multi-head attention (MHA) mechanism, using the context vector $h_{c}$ as the query and $H^{t}\in\mathbb{R}^{|F^{t}|\times d_{h}}$ as the key and value:
(12) $h^{\prime}_{c}=\text{MHA}(h_{c},W^{\textit{key}}h^{c}_{F},W^{\textit{val}}h^{c}_{F}).$
(13) $h^{\prime}_{c}=h^{\prime}_{c}+\text{IDT}(h_{c}),$
(14) $q_{c}=h^{\prime}_{c}+\text{MLP}(h^{\prime}_{c}).$
(15) $p_{i}=\text{Softmax}\left(C\cdot\tanh\left(\frac{(q_{c})^{T}W^{\textit{logit}}h^{c}_{F}}{\sqrt{d_{h}}}-\log(\text{dist}_{i})\right)\right)_{i},$
6. Experiments
Figure 4. Study of our proposed model with different initial contexts – coordinates, distances, and our Contextual Gating (CG) – as well as Neural Adaptive Bias (NAB). CG and NAB perform the best, particularly in out-of-distribution (OOD) cases.
6.1. Experimental Setup
Classical Baselines.
@@ 6.1. Experimental Setup
In the experiments, we compare three SOTA traditional optimization approaches: LKH3(Helsgaun, 2017): a heuristic algorithm with strong performance on (A)TSP problems, PyVRP(Wouda et al., 2024): a specialized solver for VRPs with comprehensive constraint handling capabilities; and Google OR-Tools(Perron and Didier, 2024): a versatile optimization library for various CO problems.
Learning-Based Methods.
@@
We compare against SOTA NCO methods divided in two categories. 1) Node-only encoding learning methods: POMO(Kwon et al., 2020), an end-to-end multi-trajectory RL-based method based on attention mechanisms; MTPOMO(Liu et al., 2024), a multi-task variant of POMO; MVMoE(Zhou et al., 2024a), a mixture-of-experts variant of MTPOMO; RF(Berto et al., 2024c): an RL-based foundation model for VRPs; ELG(Gao et al., 2024), a hybrid of local and global policies for routing problems; BQ-NCO(Drakulic et al., 2023): a decoder-only transformer trained with supervised learning; LEHD(Luo et al., 2023): a supervised learning-based heavy decoder model. 2) Node and edge encoding learning methods: GCN(Duan et al., 2020): a graph convolutional network with encoding of edge information for routing; MatNet(Kwon et al., 2021): an RL-based solver encoding edge features via matrices and finally GOAL(Drakulic et al., 2024a): a generalist agent for several CO problems, including routing problems.
Training Configuration.
@@
We perform training runs under the same settings for fair comparison for our model, MatNet for ATSP and ACVRP, and GCN for ACVRP. Node-only models do not necessitate retraining since our datasets are already normalized in the $[0,1]^{2}$ coordinates ranges (with locations sampled uniformly), and we do not retrain supervised-learning models since they would necessitate labeled data. We train with the Adam optimizer with an initial learning rate of $4\times 10^{-4}$, which decays by a factor of 0.1 at epochs 180 and 195. Training is completed within 24 hours on 4×4$\times$4 × NVIDIA A100 40GB GPUs using a batch size of 256, processing 100,000 instances per epoch. The model follows a Transformer-based architecture with 128-dimensional embeddings, 8 attention heads, 512-dimensional feedforward layers, and 12 transformer layers in total (further details in Appendix B). The training dataset consists of 80 cities with instances randomly generated from subsampled real-world city base map topologies with the remaining 20 reserved for OOD testing (see LABEL:tab:city_details for train/test splits).
Testing Protocol.
@@
The test data consists of in-distribution evaluation for 1) In-dist: new instances generated from the 80 cities seen during training, 2) OOD (city) out-of-distribution generalization over new city maps and 3) OOD (cluster) out-of-distribution generalization to new location distributions across maps. The test batch size is set to 32, and a data augmentation factor of 8 is applied to all models except supervised learning-based ones, i.e., LEHD, BQ-NCO, and GOAL. All evaluations are conducted on an NVIDIA A6000 GPU paired with an Intel(R) Xeon(R) CPU @ 2.20GHz.
6.2. Main Results
Table 1 shows the results between our and the baseline models across ATSP, ACVRP, and ACVRPTW tasks, with inference times indicated in parentheses. The results clearly demonstrate that our method achieves state-of-the-art performance across all problem settings, consistently outperforming existing neural solvers in both solution quality and computational efficiency.
@@ 6.2. Main Results
Notably, unlike previous approaches that require separate models for different problem types, our method effectively handles all routing problems within a single unified framework. This key advantage highlights the model’s adaptability and scalability across diverse problem instances while maintaining strong generalization capabilities for both in-distribution and out-of-distribution scenarios in real-world setting.
6.3. Analyses

### v2

@@ 4.1.1 Adaptive Node Embedding
The Adaptive Node Embedding module synthesizes distance-related features with node characteristics to create comprehensive node representations. A key aspect of our approach is effectively integrating two complementary spatial features: distance matrix information and coordinate-based relationships. For distance matrix information, we employ a selective sampling strategy that captures the most relevant node relationships while maintaining computational efficiency. Given a distance matrix $\mathbf{D}\in\mathbb{R}^{N\times N}$, we sample $k$ nodes for each node $i$ according to probabilities inversely proportional to their distances:
$p_{ij}=\frac{1/d_{ij}}{\sum_{j=1}^{N}1/d_{ij}}$ (3)
@@
$\mathbf{f}_{\text{dist}}=\text{Linear}(\mathbf{d}_{\text{sampled}})$ (4)
@@
$\mathbf{f}_{\text{coord}}=\text{Linear}(\mathbf{x}_{\text{coord}})$ (5)
@@
$\mathbf{h}=\mathbf{g}\odot\mathbf{f}_{\text{coord}}+(1-\mathbf{g})\odot\mathbf{f}_{\text{dist}}$ (6)
@@
$\mathbf{g}=\sigma(\text{MLP}([\mathbf{f}_{\text{coord}};\mathbf{f}_{\text{dist}}]))$ (7)
@@
$\mathbf{h}^{r}_{\text{comb}}=\text{MLP}([\mathbf{h}^{r};\mathbf{f}_{\text{node}}])$ (8)
$\mathbf{h}^{c}_{\text{comb}}=\text{MLP}([\mathbf{h}^{c};\mathbf{f}_{\text{node}}])$ (9)
@@
where $\mathbf{f}_{\text{node}}$ represents additional node features such as demand or time windows, which are transformed by an additional linear layer. This dual embedding approach allows the RRNCO model to better capture and process asymmetric relationships in real-world routing scenarios.
4.1.2 Neural Adaptive Bias for AAFM
Having established comprehensive node representations through our adaptive embedding approach, RRNCO employs an Adaptation Attention-Free Module (AAFM) based on Zhou et al. (2024a) to model complex inter-node relationships. The AAFM operates on the dual representations $\mathbf{h}^{r}_{\text{comb}}$ and $\mathbf{h}^{c}_{\text{comb}}$ to capture asymmetric routing patterns through our novel Neural Adaptive Bias (NAB) mechanism. The AAFM operation is defined as:
$\mathbf{\text{AAFM}}(Q,K,V,A)=\sigma(Q)\odot\frac{\exp(A)\cdot(\exp(K)\odot V)}{\exp(A)\cdot\exp(K)}$ (10)
where $Q=\mathbf{W}^{Q}\mathbf{h}^{r}_{\text{comb}}$, $K=\mathbf{W}^{K}\mathbf{h}^{c}_{\text{comb}}$, $V=\mathbf{W}^{V}\mathbf{h}^{c}_{\text{comb}}$, with learnable weight matrices $\mathbf{W}^{Q}$, $\mathbf{W}^{K}$, $\mathbf{W}^{V}$. While Zhou et al. (2024a) defines the adaptation bias $A$ heuristically as $\alpha\cdot\log(N)\cdot d_{ij}$ (with learnable $\alpha$, node count $N$, and distance $d_{ij}$), we introduce a Neural Adaptive Bias (NAB) that learns asymmetric relationships directly from data. NAB processes distance matrix $\mathbf{D}$, angle matrix $\mathbf{\Phi}$ with entries $\phi_{ij}=\operatorname{arctan2}(y_{j}-y_{i},x_{j}-x_{i})$, and optionally duration matrix $\mathbf{T}$, enabling joint modeling of spatial-temporal asymmetries inherent in real-world routing. Let $\mathbf{W}_{D},\mathbf{W}_{\Phi},\mathbf{W}_{T}\in\mathbb{R}^{E}$:
$\mathbf{D}_{emb}$ $=\text{ReLU}(\mathbf{D}\mathbf{W}_{D})\mathbf{W}^{\prime}_{D}$ (11)
$\bm{\Phi}_{emb}$ $=\text{ReLU}(\bm{\Phi}\mathbf{W}_{\Phi})\mathbf{W}^{\prime}_{\Phi}$ (12)
$\mathbf{T}_{emb}$ $=\text{ReLU}(\mathbf{T}\mathbf{W}_{T})\mathbf{W}^{\prime}_{T}$ (13)
We then apply contextual gating to fuse these heterogeneous information sources. When duration information is available, we employ a multi-channel gating mechanism with softmax normalization:
$\mathbf{G}=\text{softmax}\left(\frac{\left[\mathbf{D}_{\text{emb}};\ \mathbf{\Phi}_{\text{emb}};\ \mathbf{T}_{\text{emb}}\right]\mathbf{W}_{G}}{\exp(\tau)}\right)$ (14)
where $\left[\mathbf{D}_{emb};\mathbf{\Phi}_{emb};\mathbf{T}_{emb}\right]\in\mathbb{R}^{B\times N\times N\times 3E}$ is the concatenation of all embeddings, $\mathbf{W}_{G}\in\mathbb{R}^{3E\times 3}$ is a learnable weight matrix, and $\tau$ is a learnable temperature parameter. The fused representation is computed as:
$\mathbf{H}=\mathbf{G}_{1}\odot\mathbf{D}_{\text{emb}}+\mathbf{G}_{2}\odot\mathbf{\Phi}_{\text{emb}}+\mathbf{G}_{3}\odot\mathbf{T}_{\text{emb}}$ (15)
Finally, the adaptive bias matrix $\mathbf{A}$ is obtained by projecting the fused embedding $\mathbf{H}$ to a scalar value:
$\mathbf{A}=\mathbf{H}\mathbf{w}_{out}\in\mathbb{R}^{B\times N\times N}$ (16)
where $\mathbf{w}_{out}\in\mathbb{R}^{E}$ is a learnable weight vector. The resulting $\mathbf{A}$ matrix serves as a learned inductive bias that captures complex asymmetric relationships arising from the interplay between distances, directional angles, and travel durations. The Neural Adaptive Bias is then incorporated into the Adaptation Attention Free Module (AAFM). Specifically, we employ the operation defined in (10) by replacing the generic matrix $A$ with the adaptive matrix generated by NAB.
The Neural Adaptive Bias (NAB), applied through the Adaptation Attention Free Module (AAFM), yields final node representations $h^{r}_{\textit{F}}$ and $h^{c}_{\textit{F}}$ after $l$ passes through AAFM. These representations result from RRNCO’s encoding process, leveraging joint modeling of distance, angle, and duration to capture complex asymmetric patterns in real-world routing networks.
4.2 Decoder
4.2.1 Decoder Architecture
The decoder architecture integrates key elements from ReLD (Huang et al., 2025) and MatNet (Kwon et al., 2021) to autoregressively construct solutions using the dense node embeddings produced by the encoder. At each decoding step $t$, it takes as input the row and column node embeddings alongside a context vector that encapsulates the current partial solution state, such as the last visited node and dynamic attributes like remaining capacity. This context serves as the query in a multi-head attention mechanism to aggregate information from the embeddings, followed by residual connections and a multi-layer perceptron to refine the query vector. The resulting query is then employed in a compatibility layer to compute selection probabilities for feasible nodes, incorporating a negative logarithmic distance heuristic to prioritize nearby options and enhance exploration efficiency. This design enables our model to dynamically adapt static embeddings to evolving contexts, yielding strong performance across vehicle routing problems; for full technical details, including equations and implementation specifics, please refer to the Appendix A.
5 Real-World VRP Dataset
A significant challenge in applying Neural Combinatorial Optimization (NCO) to real-world routing is the lack of realistic datasets. Most existing benchmarks rely on synthetic instances with symmetric, Euclidean distances, failing to capture the complexities of actual road networks, such as one-way streets and traffic-dependent travel times, which lead to asymmetric distance and duration matrices. To bridge this gap, we introduce a new, large-scale dataset for real-world VRPs. We developed a comprehensive data generation pipeline that leverages the OpenStreetMap Routing Engine (OSRM) (OpenStreetMap contributors, 2025) to create detailed topological maps for 100 diverse cities worldwide. Each map includes location coordinates along with their corresponding asymmetric distance and duration matrices. Furthermore, we designed an efficient online subsampling method to generate a virtually unlimited number of VRP instances for training our reinforcement learning agent. This approach ensures that our model is trained on data that faithfully represents real-world routing challenges. In addition to serving as a benchmark, the dataset provides a structured basis for evaluating NCO solvers under realistic conditions and helps narrow the simulation-to-real gap, offering a useful resource for future research on practical logistics. Details of the city selection, data generation framework, and subsampling methodology are given in Appendix B.
Figure 3: Overview of our RRNCO real-world data generation and sampling framework. We generate a dataset of real-world cities with coordinates and respective distance and duration matrices obtained via OSRM. Then, we efficiently subsample instances as a set of coordinates and their matrices from the city map dataset with additional generated VRP features.
6 Experiments
@@
Table 1: Performance comparison across real-world routing tasks and distributions. We report costs and gaps calculated with respect to best-known solutions (*) from traditional solvers. Horizontal lines separate traditional solvers, node-only methods, node-and-edge methods, and our RRNCO. Lower is better $(\downarrow)$.
@@
RRNCO 122.466 3.736 52s 122.986 3.775 52s 40.937 4.290 52s
6.1 Experimental Setup
Classical Baselines
@@ 6.1 Experimental Setup
In the experiments, we compare three SOTA traditional optimization approaches: LKH3 (Helsgaun, 2017): a heuristic algorithm with strong performance on (A)TSP problems, PyVRP (Wouda et al., 2024): a specialized solver for VRPs with comprehensive constraint handling capabilities; and Google OR-Tools (Perron and Didier, 2024): a versatile optimization library for CO problems.
Learning-Based Methods
@@
We compare against SOTA NCO methods divided in two categories. 1) Node-only encoding learning methods: POMO (Kwon et al., 2020), an end-to-end multi-trajectory RL-based method based on attention mechanisms; MTPOMO (Liu et al., 2024), a multi-task variant of POMO; MVMoE (Zhou et al., 2024b), a mixture-of-experts variant of MTPOMO; RF (Berto et al., 2025b): an RL-based foundation model for VRPs; ELG (Gao et al., 2024), a hybrid of local and global policies for routing problems; BQ-NCO (Drakulic et al., 2023): a decoder-only transformer trained with supervised learning; LEHD (Luo et al., 2023): a supervised learning-based heavy decoder model and AAFM (Zhou et al., 2024a), introduced in the ICAM framework as an attention-free alternative enabling instance-conditioned adaptation. 2) Node and edge encoding learning methods: GCN (Duan et al., 2020): a graph convolutional network with encoding of edge information for routing; MatNet (Kwon et al., 2021): an RL-based solver encoding edge features via matrices; ReLD-MTL and ReLD-MoEL (Huang et al., 2025), which incorporate identity mapping and feed-forward decoder refinements to significantly improve cross-size and cross-problem generalization and GOAL (Drakulic et al., 2024): a generalist agent trained via supervised learning for several CO problems, including routing problems.
Training Configuration
@@
We perform training runs under the same settings for fair comparison for our model, MatNet for ATSP and ACVRP, and GCN for ACVRP. Node-only models do not necessitate retraining since our datasets are already normalized in the $0,1]^{2}$ coordinates ranges (with locations sampled uniformly), and we do not retrain supervised-learning models since they would necessitate labeled data. The model is trained for about 24 hours on $4\times$ NVIDIA A100 40GB GPUs, with all training settings and train, test dataset details provided in Appendix Section C.1 and Appendix Section B.4.
Testing Protocol
@@
The test data consists of in-distribution evaluation for 1) In-dist: new instances generated from the 80 cities seen during training, 2) OOD (city) out-of-distribution generalization over new city maps and 3) OOD (cluster) out-of-distribution generalization to new location distributions across maps. The test batch size is 32, and a data augmentation factor of 8 is applied to all models except supervised learning-based ones, i.e., LEHD, BQ-NCO, and GOAL. All evaluations are conducted on an NVIDIA A6000 GPU paired with an Intel(R) Xeon(R) CPU @ 2.20GHz [^1] Code: https://github.com/ai4co/real-routing-nco .
6.2 Main Results
@@
Table 1 shows the results between our and the baseline models across ATSP, ACVRP, and ACVRPTW tasks, with inference times in parentheses. The results clearly demonstrate that our method achieves state-of-the-art performance across all problem settings, consistently outperforming existing neural solvers in both solution quality and computational efficiency. Notably, unlike previous approaches that require separate models for different problem types, our method effectively handles all routing problems within a single unified framework. This key advantage highlights the model’s adaptability and scalability across diverse problem instances while maintaining strong generalization for both in-distribution and out-of-distribution scenarios in real-world settings.
6.3 Analyses
Ablation study on proposed components
We perform an ablation study on proposed model components in Fig. 4: initial contexts with coordinates, distances, and our Adaptive Node Embedding(ANE), as well as the Neural Adaptive Bias (NAB). We find ANE and NAB perform the best, particularly in out-of-distribution (OOD) cases. Remarkably, in cluster distributions, the NAB shows a relative improvement greater than $15\%$.
Figure 4: Study of our proposed model with different initial contexts: coordinates, distances, Adaptive Node Embedding (ANE), and Neural Adaptive Bias (NAB). ANE and NAB perform best, particularly in out-of-distribution (OOD) cases.
Granular ablation on NAB inputs
To further investigate the contribution of each input modality to the Neural Adaptive Bias (NAB), we conduct a granular ablation study on ACVRPTW in Table 2. Starting from the full model that jointly uses distance, duration, and angle matrices, we progressively remove components. The results show that each modality contributes to the final performance: removing duration increases the gap from 3.74% to 3.93% (in-distribution), and further removing angle increases it to 4.06%. The performance degradation is more pronounced in the OOD (cluster) setting, where the gap increases from 4.30% to 4.84%, demonstrating that the joint modeling of all three features is essential for robust generalization.
Table 2: Granular ablation on NAB inputs for ACVRPTW. We progressively remove duration and angle from the full NAB model.
Model In-distribution OOD (city) OOD (cluster)
Cost Gap (%) Cost Gap (%) Cost Gap (%)
PyVRP 118.056 * 118.513 * 39.253 *
RRNCO Full (D + T + $\Phi$) 122.467 3.74 122.986 3.78 40.937 4.29
- Duration (D + $\Phi$) 122.693 3.93 123.249 4.00 41.077 4.65
- Duration - Angle (D only) 122.849 4.06 123.364 4.09 41.151 4.84
Importance of real-world data generators